# Federated Cubic Regularized Newton Learning with Sparsification-amplified Differential Privacy

### v1

@@ Appendix C Proof of Lemma 3
$O\left(\left(\frac{L^{2}}{\mu\tau}+\frac{kT\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\right)\right).$
Then, by setting the number of local iterations to $\tau=\frac{L^{2}\varepsilon^{2}m^{2}}{kT\log(1/\delta_{0})(G_{1}+G_{2}D)^{2}}$, we obtain that the subotimality is given by (10).
@@ Appendix D Proof of Theorem 2
$\leq$ $O\left(\frac{k\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}T}{\varepsilon^{2}m^{2}\mu}\right)\triangleq\Gamma_{1}.$
@@
For the convergence of Phase I, inspired by Nesterov and Polyak [33], we let $\tilde{u}_{t+1}=\frac{9}{4}u_{t}$, and assume $u_{t}\geq\frac{3\Gamma_{1}}{\lambda}$. Then, there is $\tilde{u}_{t+1}\geq 1$ and the evolution of Phase I becomes:
@@
$=$ $O\left(\frac{k\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\cdot T\right),$

### v2

@@ Appendix C Proof of Lemma 2
$O\left(\left(\frac{L^{2}}{\mu\tau}+\frac{kT\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\right)\right).$
Then, by setting the number of local iterations to $\tau=\frac{L^{2}\varepsilon^{2}m^{2}}{kT\log(1/\delta_{0})({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)^{2}}$, we obtain that the subotimality is given by (11).
@@ Appendix D Proof of Theorem 2
$\leq$ $O\left(\frac{k\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}T}{\varepsilon^{2}m^{2}\mu}\right)\triangleq\Gamma_{1}.$
@@
For the convergence of Phase I, inspired by Nesterov & Polyak (2006), we let $\tilde{u}_{t+1}=\frac{9}{4}u_{t}$, and assume $u_{t}\geq\frac{3\Gamma_{1}}{\lambda}$. Then, there is $\tilde{u}_{t+1}\geq 1$ and the evolution of Phase I becomes:
@@
$=$ $O\left(\frac{k\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\cdot T\right),$